OpenStudy (anonymous):

Find the equation of the straight line that: passes through (2, 3) and has a gradient 3 using y = mx + c. Anyone can explain please?

OpenStudy (anonymous):

y = 3x + c 3 = 3(2) + c 3 = 6 + c -3 = c Therefore y = 3x -3

OpenStudy (anonymous):

Thank you Omega, appreciate it!